# Black-box Membership Inference Attacks against Fine-tuned Diffusion Models

### v1

@@ 2.4.2 Black-box MIA against Recent Diffusion Models
Wu et al. [61] leveraged the intuition that the generated samples exhibit a higher degree of fidelity in replicating the training samples, and demonstrate greater alignment with their accompanying textual description. However, the authors only tested their attack on off-the-shelf models with explicitly known training sets. In the realistic setting where the training set is unknown (which is the purpose of MIAs), their attack cannot work.
@@
Additionally, Zhang et al. [63] trained a classifier based on samples generated by the target model (labelled 1) and samples not used in training (labelled 0). The classifier can then determine whether the target sample was used in training. However, it needs to (1) know the non-training samples, and (2) ensure the two distributions (of generated samples and non-training samples) are different enough. Both conditions are not necessarily true in a realistic setting.
3 Methodology
@@
Drawing inspiration from GAN-Leaks, our approach aims to determine the membership of query samples through similarity scores analysis. However, GAN-Leaks relies on Parzen window density estimation [38] to estimate the probability of model-generated query samples, which often results in non-stable probability estimates. We propose utilizing the intrinsic characteristics of diffusion models with formal proofs. Specifically, we leverage the training objective of diffusion models to more directly and intuitively represent the use of similarity scores as an attack feature for determining the membership of query samples.
Algorithm 1 High-level Overview of Our Attack.
1:Target sample $x$, target model $\mathcal{M}$, distance metrics $S(\cdot,\cdot)$, the image captioning model $\mathcal{C}$, and the image feature extractor $E$.
2:if $T_{q}$notin $x$ then
3:     $T_{q}$ = $\mathcal{C}(I_{q})$triangleright Synthesize the text for $\mathcal{M}$.
4:end if
5:for $i=1$ to $m$ do $\triangleright$ Perform $m$ repetitive queries.
6:     $I_{g}^{i}$ = $\mathcal{M}(T_{q})$
7:end for
8:$\mathcal{A}\left[\langle S(E(I_{g}),E(I_{q}^{i}))\rangle_{i=1}^{m}\right]$$\triangleright$ MIA results.
Moreover, our attack is tailored for modern diffusion models, which (compared to GANs and VAEs) have the capability of taking a text prompt and original image as input. Specifically, we utilize an image feature extractor and a captioning model, as detailed below.
@@
As previously mentioned in Section 2.3.1, in some scenarios, the data record $x$ lacks the text component. Consequently, we resort to a captioning model to generate the corresponding text. For our experiments, we utilize BLIP2 [27] as the captioning model. To ensure that the generated textual descriptions closely match the style of the model’s training dataset, we also consider further use of query data and the auxiliary dataset to fine-tune the captioning model.
@@
Algorithm 1 gives the high-level overview of our attack. The intuition is to compare the generated images with the target image and compute a similarity score used for MIAs (specific instantiations of $\mathcal{A}$ to be presented in Section 3.2). According to whether the text is available or not, we might need the captioning model to synthesize the text. Once the captioning is complete, we repeatedly query the victim model $m$ times to generate $m$ images. For each generated image, we calculate a similarity score relative to the query image. Finally, we return these as an $m$-dimensional similarity score to determine the target/query data’s membership.
@@
Note that while the attack pipeline is perhaps straightforward, its intuition relies on the formal analysis of the diffusion models. In what follows, we first describe its theoretical foundation, and then instantiate it with different MIA paradigms based on the output score.
3.1 Theoretical Foundation
We aim to establish a theory demonstrating the distance between the query image $I_{q}$ and generated image $I_{g}$ can be used as a metric to infer the membership of $x$. We leverage the internal property of the diffusion model, which is inherently structured to optimize the log-likelihood: If $x$ is in the training set, its likelihood of being generated should be higher. However, due to the intractability of calculating log-likelihood in diffusion models, these models are designed to use the Evidence Lower Bound (ELBO) as an approximation of log-likelihood [18], as shown later in Equation 3. In Theorem 1, we first argue that ELBO of the diffusion model can be interpreted as a chain of generating images at any given timestep that approximates samples in the training set. Then, in Theorem 2, based on the loss function of the Stable Diffusion [44], we extend the result and demonstrate that this argument remains valid. Therefore, we can reasonably employ the similarity between the generated images and the query image as our attack. Note that GAN-Leaks [7] also shares this intuition of using similarity. However, it relies more on intuition and lacks a solid foundation, as the training of GANs is different (not a streamlined process as in diffusion models).
We elaborate on the theoretical foundation of our attack in the following detailed discussion.
Theorem 1.
Assuming we have a well-trained diffusion model $\hat{x}_{\theta}$ [^7] We previously use $\mathcal{U}_{\theta}$ to denote U-Net, now by slightly abusing notations we use $\hat{x}_{\theta}$ for easier presentations., the high similarity scores between the query data $x$ and its generated image $\hat{x}_{\theta}(x_{t},t)$, the higher the probability of $\mathsf{Pr}\left[x\in\mathcal{D}_{m}|x,\theta\right]$.
$\mathsf{Pr}\left[x\in\mathcal{D}_{m}|x,\theta\right]\propto-{\left\|x_{0}-\hat{x}_{\theta}(x_{t},t)\right\|^{2}_{2}}$
where $\theta$ denotes the parameters of the model.
Proof.
Diffusion models employ the ELBO to approximate the log-likelihood $p(x)$ of the entire training dataset.
$\cdots$
The primary focus of optimization is on $L_{t-1}$, as explicated in the original work [18]. The other terms are treated as constants and an independent decoders. Approximately, the objective function can be represented as
$\min\mathbf{D}_{KL}(q(x_{t-1}|x_{t},x_{0})\;\|\;p_{\theta}(x_{t-1}|x_{t})).$
Based on the assumption in DDPM [18], to elucidate further:
[…]
$\mathsf{Pr}\left[x\in\mathcal{D}_{m}|x,\theta\right]\propto-{\left\|x_{0}-\hat{x}_{\theta}(x_{t},t)\right\|^{2}_{2}}$ (6)
We have linked the probability of $x$ being a member to its similarity score with generated images in the unconditional diffusion model. Our next step is to prove this link persists in Stable Diffusion model.
Theorem 2.
[…]
$\mathsf{Pr}\left[x\in\mathcal{D}_{m}|x,\theta\right]\propto-{\left\|x_{0}-\hat{x}_{\theta}(z_{t},t,\phi_{\theta}(p))\right\|^{2}_{2}}$
Proof.
In the original paper [44], the loss function of the Stable Diffusion model is described as follows:
The latent code $z_{t}$ is of a much smaller dimension than that of the original image. The denoising network $\mathcal{U}_{\theta}$ predicts the noise at timestep $t$ based on $z_{t}$ and the embedding generated by $\phi_{\theta}$, which takes $p$ as its input. Given that the forward process of the Stable Diffusion [44] is fixed, Equation 1 remains applicable. Therefore, by substituting in the expression $\epsilon_{t}=\frac{x_{t}-\sqrt{\bar{\alpha_{t}}}x_{0}}{\sqrt{1-\bar{\alpha}_{t}}}$ and discarding other weight terms, we can rederive the loss function of the Stable Diffusion model as:
$L_{LDM}=\mathbb{E}_{\mathcal{E}(x),t}\left[\|x_{0}-\hat{x}_{\theta}(z_{t},t,\phi_{\theta}(p))\|^{2}_{2}\right]$ (7)
As seen from Equation 7, Stable Diffusion is essentially trained to optimize image predictions at any given timestep to closely approximate the original $x_{0}$. For the Stable Diffusion model, we can still distinguish between member samples and non-member samples by the similarity scores $||x_{0}-\hat{x}_{\theta}(z_{t},t,\phi_{\theta}(p))||^{2}_{2}$, which is expressed as:
$\mathsf{Pr}\left[x\in\mathcal{D}_{m}|x,\theta\right]\propto-{\left\|x_{0}-\hat{x}_{\theta}(z_{t},t,\phi_{\theta}(p))\right\|^{2}_{2}}$ (8)
In Equation 8, we prove that under the attack situation, when the model has pre-trained parameters $\theta$, the probability of $x$ being a member set sample is inversely proportional to
$\left\|x_{0}-\hat{x}_{\theta}(x_{t},t,\phi_{\theta}(p))\right\|^{2}_{2}.$
Since this representation quantifies the distance between the predicted original image and the actual ground truth during training, we can denote the image generated by model as $I_{g}$ (which also corresponds to $\hat{x}$ in Section 2.1.2). The similarity score between $I_{g}$ and $I_{q}$ from the query data $x$ can be represented as $S(I_{q},I_{g})$. Here, $S$ is a distance metric (i.e., Cosine similarity, $\ell_{1}$ or $\ell_{2}$ distance, or Hamming distance). Given that a higher similarity score indicates a higher probability of the data being a training sample, inference model can be formulated accordingly.
3.2 Instantiations
@@
Utilizing the scores obtained from Algorithm 1, we instantiate three different types of MIAs according to Section 2.2. In our evaluation, we try all three of them, and observe the last one is usually the most effective one.
@@
The threshold-based MIA takes the similarity scores, apply a statistical function $f$ to it, and then decide membership if the result is above a pre-defined threshold $\tau$, i.e.,
@@
Following the work by Carlini et al. [5], we know we can also use the likelihood ratio attack against diffusion models. In our analysis, we leverage similarity scores derived from shadow models to delineate two distinct distributions: $\mathbb{Q}_{in}$ and $\mathbb{Q}_{out}$. Specifically: For $\mathbb{Q}_{in}$, consider query image $I_{q}$ that belong to the member set $\mathbf{D}_{m}$. We then define $\mathbb{Q}_{in}$ as
$\mathbb{Q}_{in}$ =$=\left\{f\left[\left\langle S\left(E(I_{q}),E(I_{g}^{i})\right)\right\rangle_{i=1}^{m}\right]\;\bigg{|}\;I_{q}\in\mathbf{D}_{m}\right\}.$
Similarly, for $\mathbb{Q}_{out}$, when query image $I_{q}$ are part of the non-member set $\mathbf{D}_{nm}$, we have
$\mathbb{Q}_{out}$ =$=\left\{f\left[\left\langle S\left(E(I_{q}),E(I_{g}^{i})\right)\right\rangle_{i=1}^{m}\right]\;\bigg{|}\;I_{q}\in\mathbf{D}_{nm}\right\}.$
For target query point $x$, membership inference can be deduced by assessing:
@@
Given that the obtained similarity score is represented as a high dimensional vector, the classifier-based MIA feeds $f\left[\left\langle S\left(E(I_{q}),E(I_{g})\right)\right\rangle_{i=1}^{m}\right]$ directly into a classifier (we use a multilayer perceptron in our evaluation). This approach aligns with the methods of Shokri et al. [51], leveraging the machine learning model as the inference model to execute the attack.
@@
4 Experiment Setup
4.1 Datasets
In this study, we primarily employ the MS COCO [30], CelebA-Dialog [23], and WIT datasets [57] for evaluation. MS COCO is widely used for training and testing various text-to-image models. The CelebA-Dialog dataset, with its extensive facial data and descriptions, aligns well with our interest in sensitive data descriptions. Meanwhile, the WIT dataset, curated from web scraping Wikipedia for images and their associated text descriptions, offers a diverse range of images and distinct textual styles, serving as an excellent benchmark for assessing model robustness.
@@ 4.1 Datasets
4.2 Evaluation Metrics
To systematically evaluate the efficacy of our proposed attack, we opted for multiple evaluation metrics as performance indicators. Similar to other comparable attacks [12, 20, 34, 25, 61, 6, 5], we will employ Attack Success Rate (ASR), Area Under the Curve (AUC), and True Positive Rate (TPR) at low False Positive Rate (FPR) as our evaluation metrics.
We opted to use Stable Diffusion v1111-5555999https://huggingface.co/runwayml/stable-diffusion-v1-5 checkpoints as our pre-trained models. All experiments were carried out using two Nvidia A100 GPUs, and each fine-tuning of the model required an average of three days.
4.3 Competitors
For our evaluation, we first compare our work with existing black-box attacks on diffusion models [34, 14]. For our attack, based on the categorization provided in Section 2.3.1, the attacker will obtain information of two distinct dimensions, leading to four different scenarios. We call them Attack-I to Attack-IV. Below we introduce them in more detail.

### v4

@@ II-E2 Black-box MIA against Recent Diffusion Models
Wu et al. [61] leveraged the intuition that the generated samples exhibit a higher degree of fidelity in replicating the training samples, and demonstrate greater alignment with their accompanying textual description. However, the authors did not use the shadow model technique and only tested their attack on off-the-shelf models with explicitly known training sets. In the realistic setting where the training set is unknown (which is the purpose of MIAs), their attack cannot work.
@@
Additionally, Zhang et al. [63] trained a classifier based on samples generated by the target model (labeled 1) and samples not used in training (labeled 0). The classifier can then determine whether the target sample was used in training. However, it needs to (1) know the non-training samples, and (2) ensure the two distributions (of generated samples and non-training samples) are different enough. Both conditions are not necessarily true in a realistic setting.
III Methodology
@@
Drawing inspiration from GAN-Leaks, our approach aims to determine the membership of query samples through similarity score analysis. However, GAN-Leaks relies on Parzen window density estimation to estimate the probability of query samples [38]. This method often results in unstable probability estimates due to the large sampling size, as we mentioned in Section II-E1. We propose utilizing the intrinsic characteristics of diffusion models with formal proofs to design a more efficient and suitable attack for diffusion models. Specifically, we leverage the training objective of diffusion models to more directly and intuitively represent the use of similarity scores as an attack feature for determining the membership of query samples.
III-A Theoretical Foundation
In this section, we aim to establish a detailed theory demonstrating the distance between the query image $I_{q}$ and generated image $I_{g}$ can be used as a metric to infer the membership of $x$. We leverage the internal property of the diffusion model, which is inherently structured to optimize the log-likelihood: If $x$ is in the training set, its likelihood of being generated should be higher. However, due to the intractability of calculating log-likelihood in diffusion models, these models are designed to use the Evidence Lower Bound (ELBO) as an approximation of log-likelihood [18], as shown later in Equation 7. In Theorem 1, we first argue that ELBO of the diffusion model can be interpreted as a chain of generating images at any given timestep that approximates samples in the training set. Then, in Theorem 2, based on the loss function of the Stable Diffusion [44], we extend the result and demonstrate that this argument remains valid. Therefore, we can reasonably employ the similarity between the generated images and the query image as our attack. Note that GAN-Leaks [7] also shares this intuition of using similarity. However, it relies more on intuition and lacks a solid foundation, as the training of GANs is different (not a streamlined process as in diffusion).
From the perspective of the training process, we proposed these two theorems that facilitate our attack.
Theorem 1.
Assuming we have a pre-trained diffusion model $\hat{x}_{\theta}$ [^7] We previously use $\mathcal{U}_{\theta}$ to denote U-Net, now by slightly abusing notations we use $\hat{x}_{\theta}$ for easier presentations. with its training set $\mathcal{D}_{m}$, and use a bit $b$ to represent the membership of query sample $x$ ($1$ for member and 00 for non-member). The higher similarity scores between the query data $x$ and its generated image $\hat{x}_{\theta}(x_{t},t)$, the higher the probability of $\mathsf{Pr}\left[b=1|x,\theta\right]$.
$\mathsf{Pr}\left[b=1|x,\theta\right]\propto-{\left\|x_{0}-\hat{x}_{\theta}(x_{t},t)\right\|^{2}_{2}}$
where $\theta$ denotes the parameters of the model.
The proof of Theorem 1 can be found at Appendix C-A. In Theorem 1, we show that the training objective for the diffusion model can be seen as reconstructing the training image at each timestep. During the training phase, the diffusion model is trained to predict the $\epsilon_{t}$ using Equation 2. Because the existing of Equation 1 and $\bar{\alpha}_{t}$ is a pre-defined hyperparameter, $\hat{x}$ can be easily calculated using $\epsilon_{t}$ at each step. Thus, a data sample from the diffusion models’ member set should have higher similarity with its replication $\hat{x}_{\theta}(x_{t},t)$ at each timestep. Naturally, the denoised output from the diffusion model should have higher similarity scores with member set samples. Our work is focused on the black-box attack; the intermediate-generated images will not be used in our attack.
In the above, we have linked the probability of query sample $x$ belonging to the member set to its similarity score with generated images in the unconditional diffusion model. For this type of diffusion model, although we can prove the training image has this property with its replica. We still cannot design the black-box attack on it because the inference process is random. We cannot control the unconditional diffusion model to reconstruct the specific data sample. This generation process is the same with VAE and GAN. Hence, the existing black box attacks are to sample a large number of images from the models [34]. And then do the Monte Carlo [17] or GAN-Leaks [7] attack.
However, we can employ this property to execute the membership inference attack with conditional diffusion models (e.g., Stable Diffusion). The main difference between Stable Diffusion and the unconditional diffusion model is that the former one can do conditional generation. According to the prompt input, Stable Diffusion can generate an image that aligns with it. Therefore, we can use prompts to guide the model and synthesize images for a specific data sample. Theorem 2 is to prove this property persists in the Stable Diffusion.
Theorem 2.
[…]
$\mathsf{Pr}\left[b=1|x,\theta\right]\propto-{\left\|\mathrm{D}(z_{0})-\mathrm{D}(\hat{z}_{\theta}(z_{t},t,\phi_{\theta}(p)))\right\|^{2}_{2}}$
Where $z_{t}$ represents the latent representation, $z_{0}=\mathcal{E}(x)$. $\mathcal{E}$ is the encoder in Stable Diffusion.
The detailed proof of Theorem 2 is presented at Appendix C-B. Compared with the unconditional diffusion model in Theorem 1, Stable Diffusion’s training objective function is similar and just adds a pre-trained encoder $\phi_{\theta}$ and transfers the diffusion process into latent space. Therefore, we can prove the probability of $x$ being a member set sample is inversely proportional to
$\left\|\mathrm{D}(z_{0})-\mathrm{D}(\hat{z}_{\theta}(z_{t},t,\phi_{\theta}(p)))\right\|^{2}_{2}.$
This representation quantifies the distance between the synthesized original image and the actual ground truth during training. Considering the realistic situation and settings, we designed four attacks (as shown in Section II-D1) to use this property in different scenarios. to use this property in different scenarios. However, for general representation, we simplify denote the image generated by the model as $I_{g}$ (which also corresponds to $\hat{x}$ in Section II-A2, $\hat{x}_{\theta}(x_{t},t)$ in Theorem 1, and $\mathrm{D}(\hat{z}_{\theta}(z_{t},t,\phi_{\theta}(p)))$ in Theorem 2). The similarity score between $I_{g}$ and $I_{q}$ from the query data $x$ can be represented as $S(I_{q},I_{g})$. Here, $S$ is a distance metric (i.e., Cosine similarity, $\ell_{1}$ or $\ell_{2}$ distance, or Hamming distance). Given that a higher similarity score indicates a higher probability of the data being a training sample, the inference model can be formulated accordingly.
III-B Attack Pipeline
According to Section III-A, our attack needs to calculate the similarity between query image $I_{q}$ and generated image $I_{g}$. We choose to compute the image embedding similarity scores by using image feature extractors. Also, to execute our attack in Attack-II and Attack-IV, we incorporate the captioning model in our work.
@@ III-B Attack Pipeline
As previously mentioned in Section II-D1, in some scenarios, the data record $x$ lacks the text component. Consequently, we resort to a captioning model to generate the corresponding text. For our experiments, we utilize BLIP2 [27] as the captioning model. To ensure that the generated textual descriptions closely match the style of the model’s training dataset, we also consider further use of the auxiliary dataset to fine-tune the captioning model.
Algorithm 1 High-level Overview of Our Attack.
1:Target sample $x$, target model $\mathcal{G}$, distance metrics $S(\cdot,\cdot)$, the image captioning model $\mathcal{C}$, the instantiation of attack $\mathcal{A}$, and the image feature extractor $E$.
2:if $T_{q}$notin $x$ then
3:     $T_{q}$ = $\mathcal{C}(I_{q})$triangleright Synthesize the text for $\mathcal{G}$.
4:end if
5:for $i=1$ to $m$ do $\triangleright$ Perform $m$ repetitive queries.
6:     $I_{g}^{i}$ = $\mathcal{G}(T_{q})$
7:end for
8:$\mathcal{A}\left[\langle S(E(I_{q}),E(I_{g}^{i}))\rangle_{i=1}^{m}\right]$$\triangleright$ MIA results.
@@
Algorithm 1 gives the high-level overview of our attack. The intuition is to compare the generated images with the target image and compute a similarity score used for MIAs (specific instantiations of $\mathcal{A}$ to be presented in Section III-C). According to whether the text is available or not, we might need the captioning model to synthesize the text. Once the captioning is complete, we repeatedly query the victim model $m$ times to generate $m$ images. For each generated image, we calculate a similarity score relative to the query image. Finally, we return these as an $m$-dimensional similarity score to determine the target/query data’s membership.
@@
Note that while the attack pipeline is perhaps straightforward, its intuition relies on the formal analysis of the diffusion models. We first describe its theoretical foundation and then instantiate it with different MIA paradigms based on the output score in the following.
III-C Instantiations
@@
Utilizing the scores obtained from Algorithm 1, we instantiate three different types of MIAs according to Section II-C. In our evaluation, we try all three of them, and observe the last one is usually the most effective one.
@@
The threshold-based MIA takes the similarity scores, apply a statistical function (e.g., mean, median) $f$ to it, and then decide membership if the result is above a pre-defined threshold $\tau$, i.e.,
@@
Following the work by Carlini et al. [5], we know we can also use the likelihood ratio attack against diffusion models. In our analysis, we leverage similarity scores derived from shadow models to delineate two distinct distributions: $\mathbb{Q}_{in}$ and $\mathbb{Q}_{out}$. Specifically: For $\mathbb{Q}_{in}$, consider image $I$ that belong to the member set $\mathbf{D}_{m}$. We then define $\mathbb{Q}_{in}$ as
$\mathbb{Q}_{in}$ =$=\left\{f\left[\left\langle S\left(E(I),E(I_{g}^{i})\right)\right\rangle_{i=1}^{m}\right]\;\bigg{|}\;I\in\mathbf{D}_{m}\right\}.$
Similarly, for $\mathbb{Q}_{out}$, when image $I$ are part of the non-member set $\mathbf{D}_{nm}$, we have
$\mathbb{Q}_{out}$ =$=\left\{f\left[\left\langle S\left(E(I),E(I_{g}^{i})\right)\right\rangle_{i=1}^{m}\right]\;\bigg{|}\;I\in\mathbf{D}_{nm}\right\}.$
For target query point $I_{q}$, membership inference can be deduced by assessing:
@@
Given that the obtained similarity score is represented as a high dimensional vector, the classifier-based MIA feeds $f\left[\left\langle S\left(E(I_{q}),E(I_{g}^{i})\right)\right\rangle_{i=1}^{m}\right]$ directly into a classifier (we use a multilayer perceptron in our evaluation). This approach aligns with the methods of Shokri et al. [51], leveraging the machine learning model as the inference model to execute the attack.
@@
IV Experiment Setup
IV-A Datasets
Stable Diffusion v1-5 is pre-trained on LAION-2B [48] and LAION-Aesthetics. To guarantee the integrity and effectiveness of our work, we utilize the MS COCO [30], CelebA-Dialog [23], and WIT datasets [57] for evaluation, ensuring that there is no overlap with the pretraining dataset. MS COCO is widely used for training and testing various text-to-image models. The CelebA-Dialog dataset, with its extensive facial data and descriptions, aligns well with our interest in sensitive data descriptions. Meanwhile, the WIT dataset, curated from web scraping Wikipedia for images and their associated text descriptions, offers a diverse range of images and distinct textual styles, serving as an excellent benchmark for assessing model robustness.
@@ IV-A Datasets
TABLE II: The default parameters used in Section V.
Parameters Experiment setting for our work
Training data size $100$
Epoch number $500$
Resolution $512\times 512$
Batch size $4$
Learning rate $5\times 10^{-5}$
Gradient accumulation steps $4$
Inference step $30$
Image feature extractor DeiT
Captioning model BLIP2
Distance metrics Cosine similarity
Attack type Classifier-based
IV-B Evaluation Metrics
To systematically evaluate the efficacy of our proposed attack, we opted for multiple evaluation metrics as performance indicators. Similar to other comparable attacks [12, 20, 34, 25, 61, 6, 5], we will employ Attack Success Rate (ASR), Area Under the Curve (AUC), and True Positive Rate (TPR) at low False Positive Rate (FPR) as our evaluation metrics. In Section V, all experiments are evaluated under the condition that the member set and non-member set have the same size.
We opted to use Stable Diffusion v1111-5555101010https://huggingface.co/runwayml/stable-diffusion-v1-5 checkpoints as our pre-trained models. All experiments were carried out using two Nvidia A100 GPUs, and each fine-tuning of the model required an average of three days. We presented the default fine-tuning and attack settings in Table II.
IV-C Competitors
For our evaluation, we first compare our work with existing black-box attacks on diffusion models [34, 14]. For our attack, based on the categorization provided in Section II-D1, the attacker will obtain information of two distinct dimensions, leading to four different scenarios. We call them Attack-I to Attack-IV. Below we introduce them in more detail.